Prob. 2 (50): A vehicle has a total weight of 3800 bs and the front to rear weight distribution ratio is 49/51. The wheel basFR Ideal Proportioning Curve В Split Point A FF

Could you please help with this? I will rate with a thumbs up.
Answer #1

☺ Answer. wa 3800 lbs PS: 300 psi WF WR 51. llo in Icha kin=h TR = llin=8 = 49 RTC: 0:45.mt Kp=0.63 a) ABG= 3 ( Grit Gt) b) Fporque in front azele. = Tt = ftr X Radius ftp = ME xwt = 0.45X 1862 = 837.9 lbs = fef Tf = 837.9 lbs all in = 9216.9 lbs.in=6 =2444.413 ebas (P3 ME FRE 2374.05 -0.9712 wf 2444.413 Ws = tor Static hope =1938 - 582.413 ) cor= 1355.587.0b5 Mr Far wr 22as FR = atb (FF) FER = a + b ftf of faer = utbl fret) 2202.05 atb (2374 .05) atb (837.9) 2 372,1 1829.95 - bc1536.15) =) b =please upvote